-4x(x+3)+9x=3(x+4)
Answer by stanbon(75887)

-4x(x+3)+9x=3(x+4)
-4x^2-12x+9x = 3x+12
-4x^2-3x = 3x+12
4x^2+6x+12 = 0
2(2x^2+3x+6) = 0
2x^2+3x+6 = 0
---
Solve using the quadratic formula.
------
Cheers,
Stan H.
==========